OpenStudy (anonymous):

change this fraction o a percent 2/9

OpenStudy (anonymous):

remember how to change a fraction into a decimal?

OpenStudy (anonymous):

xs it by 100 idk im on a timed test n this sucks

OpenStudy (anonymous):

like now?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

k, why dont you ask lig or saliback?

OpenStudy (anonymous):

idk they may think im stupid idk. i get all a's in my classes but this math stuff... ugh forget it

OpenStudy (anonymous):

I just wanna follow my dream n be a human service worker...i wanna help people n if i can't pass this idk it wont happen

OpenStudy (anonymous):

:( well after your test heres a video: sooo funny! tell me how it was after you watch the video. http://www.youtube.com/watch?v=owK5tHjL0aE remember, just try your best a work hard!

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)

OpenStudy (slaaibak):

hmmm, to be honest, I forgot how to do this. I know most of the fractions by heart. This one is 22.222%

OpenStudy (anonymous):

hahahahahaha awesome :))))))))))))))))))))

OpenStudy (slaaibak):

Maybe try division: 9/2 Nine does not go into 2, so then it's 0.xxxx Nine does go into twenty 2 times, so it's 2 remainder 2 0.2xxxxx Now again, nine goes into 20, so it's 2 remainder 2: 0.22xxxxx This process goes on and on and one so the fraction = 0.22222222 etc Then x 100 to get the percentage 22.222%

OpenStudy (anonymous):

so its 22 2/9 or 2 2/9

OpenStudy (anonymous):

oh wow u guys are smart haha

OpenStudy (slaaibak):

jen, there's no such thing as stupid people :) so nope, I don't think you are stupid

OpenStudy (anonymous):

well thanks but i can't get this math stuff. it's liek i totally blank out

OpenStudy (slaaibak):

It's \[{22 } {2 \over 9} %\]

OpenStudy (slaaibak):

%

OpenStudy (anonymous):

hahahahahaha you not stupid just beautiful ;)))))))))))))

OpenStudy (anonymous):

...

OpenStudy (slaaibak):

Jen, if I can give you a tip, study fractions by heart. Know that: 1/1 = 1 1/2 = 0.5 1/3 = 0.333 1/4 = 0.25 1/5 = 0.2 1/6 = 0.1666666 1/7 = 0.14285 1/8 = 0.125 1/9 = 0.111111 1/10 = 0.1 1/11 = 0.09090909 1/12 = 0.08333 I don't think you should need to know further than that

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

1/10=__ %

OpenStudy (slaaibak):

Makes math and everyday life so much easier. 1/10 = 0.1 Therefore 0.1 * 100 = 10%

OpenStudy (anonymous):

if u guys can just give me the answers so i can get through this test would be awesome!!!

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

65% of 228 is?

OpenStudy (anonymous):

good luck :)

OpenStudy (anonymous):

u guys r awesome thanks for helping me?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

37.5% of 10 is

OpenStudy (anonymous):

65% of 228 = 148.2

OpenStudy (anonymous):

37.5% of 10 is 3.75

OpenStudy (anonymous):

asadmilik...thank u!

OpenStudy (anonymous):

it`s Asad Malik not asadmilik

OpenStudy (anonymous):

sorry computer froze :(

OpenStudy (anonymous):

hahahahaha np ;)

OpenStudy (anonymous):

50% of ___is 15

OpenStudy (anonymous):

30

OpenStudy (anonymous):

ty......90% of what number is $54.00

OpenStudy (anonymous):

60

OpenStudy (anonymous):

__% of 50 is 140

OpenStudy (anonymous):

$48 is __% of 50

OpenStudy (anonymous):

35.71 96

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

interest is compounded annually. Find the compound amount after given time. Principle is $200.00 rate of interest 57% time 3 years. what is the compound amount rounded to the nearest cent

OpenStudy (anonymous):

i don`t know about this sorry :)

OpenStudy (anonymous):

its ok. what about this one. of the 8760 hrs in a yr one tv was on for 2190 hrs. what is the percent of this

OpenStudy (anonymous):

the amount in a savings acc increased from $250 to $270. what is the percent increase?

OpenStudy (anonymous):

a sales persons rate of commission is 5%. what is the commission from a $350.00 sale/

OpenStudy (anonymous):

17.5

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

17.5 commission

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

do u know the other ones or no?

OpenStudy (anonymous):

no :((((( sorry

OpenStudy (anonymous):

its ok u helped me so much thank u :)

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

by dividing the top dog into the house